OpenStudy (anonymous):

5 Questions! Think You Can Do Them? 1. 4.2p=8.4 What is P 2. 3/10=2x What Is X? 3. 2/5x=4/3 What Is X? 4. 11/2+y=9/2 What is Y? 5. -9.2= m/3.4 What Is M?

OpenStudy (anonymous):

p = 2 x = 3/20 x = 10/3 y = -1 m = -31.28

OpenStudy (anonymous):

Thaanks im a 6th grader and i just learned this 3 weeks ago! And i dont remeber how to do it! >.<